OpenStudy (anonymous):

this is a trapezoid and i am trying to find x 75cm^2=4x^2-5x

OpenStudy (anonymous):

Looks like you just have an equation to solve x for....well that's what I take out of it anyhow. So solve \[4x^2-5x-75=0\] for x.

OpenStudy (anonymous):

ok here is what they gave me bottom base is 3x-7 right side is 2x top base is x+2

OpenStudy (anonymous):

I am tryin to find x which I plugged into the area equation 75cm^2=.5(2x(3x-7+x+2)) 75cm^2=.5(6x^2-14x+2x^2+4x) 75cm^2=.5(8x^2-10x) 75cm^2=4x^2-5x .